# WebQuest: A Benchmark for Multimodal QA on Web Page Sequences

### v1

@@ Appendix C Prompts
In this section, we show the prompts used for evaluating the different MLLMs and we also discuss the improvement from prompt engineering.
C.1 Zero-shot prompt
You are given a sequence of screens and a question. Answer the question using only the information on the screens.
You should directly tell me your answer in the fewest words possible, and do not output any explanation or any other contents. Question:
C.2 Chain-of-thought prompt
You are given <num_screen> screenshots and a question. Your goal is to answer the question according to the screen information only.
Please follow the below steps to answer the question.
@@ C.2 Chain-of-thought prompt
Appendix D Data Collection Details
D.1 Data Collection Instructions
The data collection process was performed with human raters who were responsible for capturing the screenshots and generating the corresponding questions and answers. In each task, raters are asked to collect screenshots and QA on a specific website category like Shopping, Restaurants etc. The instructions given to the raters for Multi Screen QA are shown below:
* Consider use cases where a user is researching a product to buy or comparing options from different web-sites.
* Capture 3-5 web screenshots taken from Google Chrome (desktop) spanning websites related to each other.
* Most of these screenshots should be from different websites.
* Generate a question-answer pair focusing on arithmetic or logical reasoning.
* Ensure that answering the question needs to use information from all the screenshots.
For the Trace QA split, similar to the Multi Screen QA split, the raters are asked to focus on specific website category. They are asked to install custom Chrome Extension which enables them to record their interactions on the screen. The chrome extension is meant to be turned on when the rater is ready to collect data. They can pause, review or delete any of the data captured. After collecting the trace, the raters add questions and answers related to the screenshots in the trace. The specific instructions are shown below:
* Consider use cases where a user is researching a product to buy or comparing options from different web-sites.
* Go to the first web page and start the trace recording.
* Navigate through the different webpages as part of the trace. Once the navigation is complete, click "stop" on the recording and review the screenshots captured.
* Delete any screenshots containing any personal information.
* Generate a question-answer pair focusing on arithmetic or logical reasoning.
* Ensure that answering the question needs to use information from about 3-5 screenshots.
D.2 Rater Details
For the Multi Screen QA task, we used 10 raters with ages between 25 and 35 based in India. The raters were employed as contractors to help with various data collection tasks apart from this task.

### v2

@@ Appendix C Prompts
In this section, we show the prompts used for evaluating the different MLLMs. These prompts correspond to the results presented in Table 1.
C.1. Zero-shot prompt
This prompt is used for evaluating the models by directly predicting the final answer. The prompt is as follows:
You are given a sequence of screens and a question. Answer the question using only the information on the screens. You should directly tell me your answer in the fewest words possible, and do not output any explanation or any other contents. Question: <question>
C.2. Chain-of-thought prompt
This prompt encourages the models to retrieve the relevant information and then combine it to arrive at the final answer. The prompt is as follows:
You are given <num_screen> screenshots and a question. Your goal is to answer the question according to the screen information only. Please follow the below steps to answer the question.